OpenStudy (anonymous):

A certain radioactive element decays over time according to the equation y=A(1/100)^t/300 where A= the number of grams present initially and t=time in years. If 1000 grams were present initially, how many grams will remain after 900 years?

OpenStudy (zehanz):

Set A=1000 and t=900 in the formula:\[y=1000\cdot \left( \frac{ 1 }{ 100 } \right)^{\frac{ 900 }{ 300 }}=1000\cdot \left( \frac{ 1 }{ 100 } \right)^3=...\]

OpenStudy (anonymous):

so what would the answer be? I plugged it into my calculator, but its not one f the answer choices..

OpenStudy (anonymous):

please help(:

OpenStudy (zehanz):

Calculator not needed, because all numbers are powers of 10: \(1000\cdot(1/100)^3=10^3 \cdot(10^{-2})^3=10^3\cdot 10^{-6}=10^{-3}=0.001\). If this is not among your answer options, then are you sure the original formula is correct?